OpenStudy (anonymous):

(5-3i)-(2+4i) (4-3i)(9+2i) (squarerootof -25)^3 5i/3-i

OpenStudy (anonymous):

u answer mine ill answer urs

hero (hero):

Nobody wants to do your homework for you

OpenStudy (anonymous):

first one is combining like terms. you get \[(5-3i)-(2+4i)=5-3i-2-4i=3-7i\]

hero (hero):

except maybe satellite

OpenStudy (anonymous):

why not?

hero (hero):

lol

OpenStudy (anonymous):

first, the point of this website it to help people. second, it is not my homework.

OpenStudy (anonymous):

you multiply \[(a+bi)(c+di)=(ac-bd)+(ad+bc)i\]

OpenStudy (anonymous):

so \[(4-3i)(9+2i) =(4\times 9+3\times 2)+(4\times 2-3\times 9)i\] or \[42-19i\]

OpenStudy (anonymous):

\[\sqrt{-25}=5i\] so \[(5i)^3=5^3i^3=125(-i)=-125i\]

OpenStudy (anonymous):

and finally \[\frac{5i}{3-i}=\frac{5i}{3-i}\times \frac{3+i}{3+i}=\frac{-5+15i}{10}\]

OpenStudy (anonymous):

which you write in standard form as \[-\frac{1}{2}+\frac{3}{2}i\]

OpenStudy (anonymous):

so satellite wanna do mine

OpenStudy (anonymous):

yeah why not

OpenStudy (anonymous):

where is it?